Circular No. 8082

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003bf AND 2003bg Two additional apparent supernovae have been reported on unfiltered

CCD images, SN 2003bf by W. M. Wood-Vasey et al. (cf. IAUC 8079) from NEAT Haleakala images and SN 2003bg by R. Chassagne (cf. IAUC 7987).

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003bf Feb. 18.45 8^{h}08^{m}26.^{s}35 +12^{o}19′56′′.2 17.6 3′′.6 W, 1′′.4 S
2003bg Feb. 25.7 4 10 59.42 −31 24 50.3 15.0: 16′′.3 W, 24′′.6 S

Additional magnitudes from the respective discoverers: SN 2003bf, Feb. 3.24, [20.8; 21.35, 17.2. SN 2003bg in MCG –05-10-15, 1979 Dec. 11.5 UT, 19.5 (Digitized Sky Survey); 2002 Nov. 7.0, [18.0:; 2003 Feb. 26.6, 15.0:.

V838 MONOCEROTIS L. S. Bernstein, Spectral Sciences, Inc.; R. J. Rudy, D. K. Lynch, W. L.

Dimpfl, S. Mazuk, and C. C. Venturini, The Aerospace Corporation; R. C. Puetter, Center for Astrophysics and Space Science, University of California, San Diego; and R. B. Perry, Langley Research Center, NASA, report the presence, in a 0.8–2.5-μm spectrum of V838 Mon obtained on 2002 Dec. 19 UT at the Lick Observatory 3-m telescope (+ Aerospace Near Infrared Imaging Spectrograph), of significant absorption features due to aluminum oxide (detected previously but not identified; cf. IAUC 8016, 8036). Also visible are the previously reported absorptions due to water vapor and CO. The AlO features are the A-X (4,0) doublet, with peak absorptions at 1.225 and 1.244 μm, and the A-X (2,0) doublet at 1.650 and 1.687 μm. The other AlO A-X bands in the vibrational sequence are less evident due to masking by much stronger water features. A very deep, broad absorption at 1.05 μm is tentatively identified as a combination of the AlO A-X (6,0) transition and the vanadium-oxide (VO) A-X (0,0) feature seen in the spectra of late-type stars. Because the AlO and VO features arise directly from the ground electronic-vibration level, they can exhibit strong absorption at the atypically low temperature (≈ 600 K) indicated by the water and CO spectral absorption profiles. Other VO features are also evident, including the A-X (1,0), (2,0), and (3,0) bands, and the B-X (0,0) and (0,1) bands. In contrast, titanium oxide, which frequently dominates the spectra of late-type stars, is less apparent in these near-infrared data, with only the E-X (0,0) band at 0.84 μm appearing as an underlying contribution to the stronger VO B-X (0,1) band. Further spectroscopic observations of V838 Mon are urged, as its envelope is chemically active and evolving.

2003 February 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green